Hydrated calcium sulfate CaSO4*n H2O contains 55.8% sulfate by mass. calculate the value for "n".
Answer #1

By using molar mass ratios, we can find the % of Ca, SO4 and H2O :


(55.8 % SO4) ((40 grams Ca /mol) / (96 grams SO4 / mol) = 23.25 % Ca

55.8% SO4 + 23.25 % Ca = 79.05 % CaSO4

100 % - 79.05 % CaSO4 = 20.95 % H2O


use molar masses to find moles, knowing that %'s are % by mass:

Moles = mass % / molar mass


(79.05 g CaSO4) (1 mole CaSO4 / 136 g) = 0.58 moles CaSO4


(20.95 H2O) (1 mole H2O / 18.01 grams) = 1.16 moles of H2O



divide by the smallest number :


0.58 moles CaSO4 / 0.58 = 1 mole CaSO4


1.16 moles of H2O / 0.58 = 2 mole H2O


Hence, empirical formula is


CaSO4. 2 H2O (gypsum)

Therefore,

n = 2 moles of water